OpenStudy (anonymous):

Find the coordinates of the missing endpoint if M is the midpoint of segment DF. F(-3, -8), M(1,-2)

OpenStudy (anonymous):

@245424 do you know how to find midpoint of a line.?

OpenStudy (anonymous):

 if x1 and y1 are coordinates of D, x2 and y2 are coordinates of F and x and y are of midpoint M. then, x = x1+x2 / 2 , y = y1 + y2 / 2. value of x,y and x2,y2 given so from above formula you can calculate x1,y1 which coordinates of D.